OpenStudy (waheguru):

PLEASE DONT GIVR THE ANSWER JUST TAKE ME STEP BY STEP UNTIL I --- GET THE ANSWER A CIRCULAR GARDEN HAS A DIAMETER OF 12M BY HOW MUCH DOES TEH DIAMETER HAVE 2 BE INCREASED SO THAT THE AREA TRIPLES

OpenStudy (pottersheep):

I will attempt to help you

OpenStudy (anonymous):

radius is (Diameter)/(2) so for the area to be tripled i guess we can simply triply the radius

OpenStudy (pottersheep):

So area of circle: πr2 Diameter: 12 so Radius :6m The area we have now: π(6)^2 = 113.1m^2 area we need = 113.1m^2 x 3 = 339.2

OpenStudy (waheguru):

yea i got that so far

OpenStudy (pottersheep):

I'm thinking :P

OpenStudy (waheguru):

hmm

OpenStudy (pottersheep):

πr2 = 339.2 r^2 = 339.3/pi

OpenStudy (waheguru):

thats what i was doing jus now =)

OpenStudy (pottersheep):

r^2 = 108 (im rounding the numbers :P) r = 10.4

OpenStudy (waheguru):

YAY I ALSO GOT THAT but the asnwers say 8.8m

OpenStudy (pottersheep):

darn. are you sure you wrote the question right?

OpenStudy (waheguru):

lets plug it in and see if it is tripled

OpenStudy (waheguru):

yea im sure bro

OpenStudy (pottersheep):

OH I GET WHY. Because you have to say how much it increases by

OpenStudy (waheguru):

?

OpenStudy (pottersheep):

Radius now is 10.4 for diameter is 20.8. Before the diameter was 12.

OpenStudy (pottersheep):

*so diameter is 20.8

OpenStudy (waheguru):

and bak 2 our andwer 10.4

OpenStudy (waheguru):

oh never mind

OpenStudy (pottersheep):

whoopdedoo

OpenStudy (waheguru):

whait how dod u get it agiin sry

OpenStudy (pottersheep):

20.8 - 12 =

OpenStudy (pottersheep):

8.8 =D

OpenStudy (waheguru):

YEYYYYYYYYYYYY WE DID IT I GOT IT 2 THX BROOOOOO

OpenStudy (pottersheep):

I'm a girl but no problem=)

OpenStudy (waheguru):

OH sry

OpenStudy (waheguru):

wat grade?

OpenStudy (pottersheep):

11 advanced functions

OpenStudy (waheguru):

wow im just in grade 9

OpenStudy (waheguru):

where u from

OpenStudy (pottersheep):

Canada, hby?

OpenStudy (waheguru):

me 2

OpenStudy (waheguru):

hudson bay?

OpenStudy (pottersheep):

Grade 9 math can be kind of confusing since it's the first year of highschool and everyone learned different things from different schools . =)

OpenStudy (pottersheep):

my school's in ontario

OpenStudy (waheguru):

yea thx 4 ur help

OpenStudy (pottersheep):

no problem

OpenStudy (waheguru):

bye

OpenStudy (pottersheep):

byebye lol

OpenStudy (anonymous):

\[\frac{A_{2}}{A_{1}}=\frac{\pi r_{2}^{2}}{\pi r_{1}^{2}}\]\[3=\frac{\pi r_{2}^{2}}{\pi 6^{2}}=\frac{r_{2}^{2}}{36}\]etc.